6.) A 11.5 ml solution of HCl has 0.531 g/L of HCl in solution. Calculate the concentration, amount of moles of HCl, and pH of the solution.
1
Expert's answer
2020-05-08T14:04:45-0400

In 1L has 0,531g HCl

In 0,0115L : m(HCl)=(0,0115L*0,531g)/1L=0,006g

n(HCl)=0,006g/36,46g/mol=1,645*10-4mol

pH=-lg[H+]=-lg(1,645*10-4)=3,78